OpenStudy (anonymous):

Can some help me on this question. You'll use the Pythagorean theorem to solve it. 165 is the hypotenuse and 88 is one leg. Set up that equation and solve for the other leg!

OpenStudy (anonymous):

a=88 and c=165 a^2+b^2=c^2 88^2+b^2=165^2 solve for b

OpenStudy (anonymous):



OpenStudy (anonymous):

refer what @kmalone99 wrote.

OpenStudy (anonymous):

thank you monokerous how come i never get best response?

OpenStudy (anonymous):

i keep getting the wrong answer. i think am doing something wrong.

OpenStudy (anonymous):

@monokerous or @kmalone99 can you help me

OpenStudy (anonymous):

88^2+b^2=165^2 7744+b^2=27225 then what?

OpenStudy (anonymous):

subtract I think

OpenStudy (anonymous):

yes and solve for b

OpenStudy (anonymous):

how I solve for b

OpenStudy (anonymous):

\[a^2+b^2=c^2\] \[b^2=c^2-a^2\] \[b=\sqrt{c^2-a^2}\]